# EvoC2Rust: A Skeleton-guided Framework for Project-Level C-to-Rust Translation

### v3

@@ Abstract.
Translating legacy C codebases to Rust is increasingly demanded for building safety-critical systems. While various approaches have emerged for this task, they face inherent trade-offs: rule-based methods often struggle to satisfy code safety and idiomaticity requirements, while LLM-based methods frequently fail to generate semantically equivalent Rust code, due to the heavy dependencies of modules across the entire codebase. In this paper, we propose EvoC2Rust, an automated framework for converting complete C projects to equivalent Rust ones. EvoC2Rust employs a skeleton-guided translation strategy for project-level translation. The pipeline consists of three stages: 1) it first decomposes the C project into functional modules, employs a feature-mapping-enhanced LLM to transform definitions and macros, and generates type-checked function stubs, which form a compilable Rust skeleton; 2) it then incrementally translates functions, replacing the corresponding stub placeholders; 3) finally, it repairs compilation errors by integrating LLM and static analysis. Through evolutionary augmentation, EvoC2Rust combines the advantages of both rule-based and LLM-based solutions. Our evaluation on open-source benchmarks and six industrial projects demonstrates the superior performance of EvoC2Rust in project-level C-to-Rust translation. The results show that our approach outperforms the strongest LLM-based baseline by 17.24% in syntax accuracy and 14.32% in semantic accuracy, while also achieving a 43.59% higher code safety rate than the best rule-based tool.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: the IEEE/ACM International Conference on Software Engineering: Software Engineering in Practice; April 12–18, 2026; Rio de Janeiro, Brazilisbn: 978-1-4503-XXXX-X/2025/02
@@ 1. Introduction
The translation of legacy C code to Rust is gaining significant attention in software engineering, spurred by critical memory safety concerns (LiW25; DARPA). Such vulnerabilities constitute the most prevalent category of critical security defects in production C systems, accounting for approximately 70% of high-severity vulnerabilities in industry reports from Google and Microsoft (UrgentNeed). This has accelerated a paradigm shift toward compile-time safety guarantees through safe programming languages (WhiteHouse). In response, Rust has emerged as a leading alternative that offers memory safety by enforcing a strict ownership and borrowing model (Williams24a).
@@
Recent work proposes rule-based tools (c2rust; ZhangDYW23; HongR23), which convert C code to Rust using primarily syntax-based transformations on a terse intermediate representation of compilation. Such translation inherently generates non-idiomatic Rust code with low-level constructs and may also contain semantic inaccuracies. (LostinTranslation). Another line of work proposes LLM-based techniques (FLOURINE; LiYGCS24; vert; SpecTra) to utilize the code comprehension capability of code LLMs. Due to the lack of parallel C-Rust training data and significant linguistic differences, these approaches often fail to produce semantically equivalent Rust code (RustRepoTrans).
While automated code translation has been extensively investigated (JiaoYLQGS23; RoziereLCL20; Zhu0R22; Yang0YK0LHMJ024; TaoYGS24; dehghan2025translating), the translation of C-to-Rust projects poses special challenges:
Challenge-1: Substantial linguistic discrepancies for ensuring safety. C and Rust differ fundamentally in their safety models. C permits loose type checking, unrestricted pointer arithmetic, and manual memory management, whereas Rust enforces strict compile-time type safety, ownership-based memory guarantees, and disciplined reference usage through its borrowing rules. Existing rule-based translators often compromise safety by relying on raw pointers, unsafe blocks, or external C functions (LostinTranslation). While LLM-based approaches can generate code with improved safety, they often lack sufficient context understanding (e.g., variable scope or pointer lifetimes), potentially leading to logical errors or memory safety violations in the translated Rust code (LAC2R). As a result, neither rule-based nor LLM-based methods alone reliably meet Rust’s stringent safety requirements.
Challenge-2: Project-level code dependence. Translating complete C projects poses greater challenges than translating isolated functions because the translated Rust project must preserve cross-module dependencies and consistent project-level structure (projectc2rust; Syzygy). Code LLMs often struggle with such large-scale context (YunLGS24), leading to broken references, inconsistent APIs, or incorrect module boundaries (LostinTranslation; apirat). Consequently, existing approaches tend to perform well only on small-scale programs (e.g., $\leq$100 LoC) (EmreSDH21; FLOURINE), and fall short in real-world project settings.
@@
We evaluate EvoC2Rust using an open-source benchmark Vivo-Bench (vivo) and a self-collected dataset of six industrial projects. Experimental results demonstrate that EvoC2Rust significantly outperforms existing baselines in project-level C-to-Rust translation. In comparison with LLM-based approaches, EvoC2Rust improves the compilation pass rate by 17.24–50.84% and the line acceptance rate by 14.32–44.79%, while also achieving marginally better code safety. When compared to rule-based methods, EvoC2Rust attains a 43.59–96.79% higher code safety rate, while maintaining competitive performance in terms of compilation and line acceptance rates. Furthermore, at the module level, EvoC2Rust achieves a 92.25% compilation success rate and an 89.53% test pass rate on industrial projects.
@@
2. Approach
2.1. Project Skeleton Construction
Figure 2. An illustration of translating a simple C project to Rust.
Given a C project, EvoC2Rust first parses the source code using Tree-sitter (treesitter) to extract key syntactic elements and abstract them into a unified project metadata representation. This metadata captures structural information including filenames, include statements, macro definitions, type declarations, global identifiers, function signatures, and a mapping table between identifier declarations to source files. They are then systematically transformed into Rust equivalents via rule-based conversion. For instance, C include dependencies are transformed to Rust use imports, and identifier declarations are converted into Rust pub use statements (e.g., pub use {Rust_filename}::{identifier_name}) using the declaration-to-file table.
Using the converted metadata, EvoC2Rust constructs a Rust project skeleton—comprising Rust constructs automatically translated from C definitions, macros, and function signatures—that accurately preserves the structural organization of the original C project (see Figure 2). Each C function is assigned a type-checked dummy counterpart in Rust using its translated signature and the unimplemented!() macro. These placeholders serve as safe stand-ins, allowing the project to compile successfully even before the actual function bodies are translated.
@@ 2.2. Module Translation with Safety-preserving Mapping
Pattern Name: Variadic Argument Mapping   Motivation: Rust does not support C-style variadic functions (e.g., va_list).   Consequence: Current LLMs cannot directly translate C’s variadic features.   Solution: Define a VaList type and va_format! macro: pub type VaList<‘a> = &‘a [&‘a dyn Display]; macro_rules! va_format { … }   Example: C Code: int snprintf(char *str, size_t size, const char *format, …) { … } Rust Code: pub fn snprintf(mut buf: Ptr<u8>, size: usize, format: Ptr<u8>, va: VaList) -> i32 { let mut fmt = format.to_string(); … return length; }
@@ 3.1. Comparison Methods
We compare EvoC2Rust against three categories of C-to-Rust translation methods: rule-based (C2Rust), LLM-based (Self-Repair and Tymcrat), and hybrid techniques (C2SaferRust). We also incorporate a direct LLM prompting baseline that translates code without repository context.
Specifically, the baseline methods evaluated in our study are as follows:
  • C2Rust (c2rust): a rule-based C-to-Rust translator that converts C code to Rust via AST analysis and manually defined transformation rules.
  • C2SaferRust (C2SaferRust): a hybrid technique built on C2Rust, augmented with LLM-based post-processing and test-driven optimization to enhance the safety and idiomaticity of the translated Rust code. We follow the original configuration and use 5 optimization iterations.
  • Self-Repair (CRUST-Bench): a multi-agent self-repair framework for project-level C-to-Rust translation. It incrementally performs file translation and uses iterative compilation and testing feedback to correct errors automatically. We adopt the default setting of 3 self-repair rounds per task.
  • Tymcrat (typemigrating): a project-level translation method that improves translation quality through type migration. It generates multiple candidate Rust signatures and refines them iteratively using compiler feedback. We use the default configuration, which generates 4 candidate signatures per function.
@@
We implement all baseline methods using their officially released code, while substituting their underlying models with those used in our study to ensure a fair comparison.
Table 3. Statistics of the Datasets
[…]
C2R-Bench avl 9 836 29 29 9 121
bzp 18 2258 92 82 14 17
md5 2 324 8 20 1 37
sha256 2 280 9 11 2 12
rapidlz 7 748 28 38 10 12
cmptlz 25 3724 122 166 27 23
Total 63 8170 288 346 63 222
@@ 3.2. Datasets
We evaluate EvoC2Rust using two benchmarks: the open-source Vivo-Bench (vivo), and our newly introduced C2R-Bench, which comprises six industrial projects.
Vivo-Bench is derived from the 2025 Vivo C-to-Rust Innovation Competition and consists of 19 algorithmic C projects. These projects contain 200 functions, 95 definitions, and 29 non-function macros, with each project spanning 1–3 files (33–630 LoC per file). Two senior engineers produced the reference Rust translations with LLM assistance. Subsequently, they validated the translations by adding tests to achieve complete function coverage, resulting in a total of 113 top-level test cases.
@@
Each project in both Vivo-Bench (vivo) and C2R-Bench provides four artifacts: the original C source code, C test cases, the reference Rust translation, and corresponding Rust test cases. Table 3 summarizes the dataset statistics.
@@ 3.3. Evaluation Strategy and Metrics
  • Incremental Compilation Pass Rate (ICompRate): The proportion of translated modules that compile successfully when integrated step-wise into the project skeleton, measuring syntactic correctness  (projectc2rust).
  • Line Acceptance Rate (AccRate) (0001KLRRSSA22): This metric evaluates the fidelity of the initial translation by comparing it to the manually corrected version. Precision captures the percentage of correct lines in the initial output, while Recall measures their preservation in the final code.
  • Code Safety Rate (SafeRate) (CRustS): The proportion of generated Rust code that is memory-safe, i.e., free of unsafe functions or blocks.
@@
  • Test Pass Rate (TestRate) (CRUST-Bench): The percentage of generated modules that pass all associated unit tests.
@@ 3.4. Implementation Details
C2Rust (c2rust) 100 100 100 0
C2SaferRust (C2SaferRust) 100 100 100 60.00
@@
Self-Repair (CRUST-Bench) 84.21 92.08 89.24 87.60
Tymcrat (typemigrating) 87.35 92.11 92.44 97.88
@@
Self-Repair (CRUST-Bench) 52.63 33.12 38.36 77.57
Tymcrat (typemigrating) 81.76 76.82 74.47 87.11
@@
C2Rust (c2rust) 99.28 98.99 98.98 1.83
C2SaferRust (C2SaferRust) 99.28 97.47 97.68 48.24
@@
Self-Repair (CRUST-Bench) 49.21 72.35 13.79 83.92
Tymcrat (typemigrating) 72.02 77.82 74.94 95.92
@@
Self-Repair (CRUST-Bench) 39.40 45.38 10.65 87.15
Tymcrat (typemigrating) 70.56 66.72 65.91 82.75
@@
We implement EvoC2Rust using DeepSeek-V3 (deepseekv3) and Qwen3-32B (Qwen3-32B) as the foundation models. Translation tasks use greedy decoding with max_tokens set to 4096. We leverage Tree-sitter v0.22.3 (treesitter) for C code parsing, the standard rustc (rustc) for compilation, and the BGE-M3 model (BGE-M3) to generate embeddings and retrieve the top-10 most relevant patterns based on cosine similarity. Iteration limits are enforced to ensure practicality: 5 rounds for bracket repair and 3 rounds for LLM refinement.
We established verified ground-truth implementations to evaluate the line acceptance rate metric. For each project, we first translated the code using the target methods. The resulting Rust artifacts were then iteratively compiled and tested. Any compilation or test failures were manually repaired by three senior engineers with assistance from Claude Sonnet 4 (claude4) until all tests passed.
All experiments are conducted on a Linux server running Ubuntu 23.10, equipped with two NVIDIA 4090 GPUs and CUDA 12.4.
@@ 4.1. Main Results (RQ1)
Specifically, EvoC2Rust improves compilation success by 44.63% and line acceptance by 54.38% over Self-Repair on C2R-Bench with DeepSeek-V3 (42.95%/55.66% with Qwen3-32B). Self-Repair’s file-level processing often exceeds LLM context limits, leading to code truncation. Compared to Tymcrat, EvoC2Rust achieves a 21.82% higher compilation rate on C2R-Bench with DeepSeek-V3 (11.79% with Qwen3-32B). While Tymcrat employs diverse signature generation and context enrichment to reduce initial errors, its lack of expert-derived structured translation guidance ultimately constrains its effectiveness in complex scenarios. Direct Prompting performs poorest, with a 38.42% average compilation success and low line acceptance (39.13% precision, 40.64% recall). This is primarily due to the inherent linguistic gaps between C and Rust. The use of our skeleton framework, which helps manage project-wide context, prevents these metrics from being even lower.

### v4

@@ Abstract.
Translating legacy C codebases to Rust is increasingly demanded for building safety-critical systems. While various approaches have emerged for this task, they face inherent trade-offs: rule-based methods often struggle to satisfy code safety and idiomaticity requirements, while LLM-based methods frequently fail to generate semantically equivalent Rust code, due to the heavy dependencies of modules across the entire codebase. Recent studies have revealed that both solutions are limited to small-scale programs. In this paper, we propose EvoC2Rust, an automated framework for converting complete C projects to equivalent Rust ones. EvoC2Rust employs a skeleton-guided translation strategy for project-level translation. The pipeline consists of three stages: 1) it first decomposes the C project into functional modules, employs a feature-mapping-enhanced LLM to transform definitions and macros, and generates type-checked function stubs, which form a compilable Rust skeleton; 2) it then incrementally translates functions, replacing the corresponding stub placeholders; 3) finally, it repairs compilation errors by integrating LLM and static analysis. Through evolutionary augmentation, EvoC2Rust combines the advantages of both rule-based and LLM-based solutions. Our evaluation on open-source benchmarks and six industrial projects demonstrates the superior performance of EvoC2Rust in project-level C-to-Rust translation. The results show that our approach outperforms the strongest LLM-based baseline by 17.24% in syntax accuracy and 14.32% in semantic accuracy, while also achieving a 43.59% higher code safety rate than the best rule-based tool.
C-to-Rust Translation, Complete Project Conversion, Skeleton Guided, Feature Mapping, Large Language Models
journalyear: 2026copyright: ccconference: 2026 IEEE/ACM 48th International Conference on Software Engineering; April 12–18, 2026; Rio de Janeiro, Brazilbooktitle: 2026 IEEE/ACM 48th International Conference on Software Engineering (ICSE-SEIP ’26), April 12–18, 2026, Rio de Janeiro, Brazildoi: 10.1145/3786583.3786856isbn: 979-8-4007-2426-8/2026/04ccs: Software and its engineering Source code generationccs: Computing methodologies Machine translation
@@ 1. Introduction
The translation of legacy C code to Rust is gaining significant attention in software engineering, spurred by critical memory safety concerns (Li et al., 2025; DARPA-SN-24-89, 2024). Such vulnerabilities constitute the most prevalent category of critical security defects in production C systems, accounting for approximately 70% of high-severity vulnerabilities in industry reports from Google and Microsoft (Lord, 2023). This has accelerated a paradigm shift toward compile-time safety guarantees through safe programming languages (Office of the National Cyber Director, 2024). In response, Rust has emerged as a leading alternative that offers memory safety by enforcing a strict ownership and borrowing model (Williams, 2024).
@@
The figure illustrates the workflow of EvoC2Rust. First, a C project is decomposed into modules containing macros, definitions, function signatures, and bodies. The extracted metadata is transformed and assembled into an initial Rust skeleton. Concurrently, module-level code translation employs feature mappings to generate corresponding Rust modules, which are integrated into the skeleton to form an initial Rust project. This project then undergoes a post-generation repair stage consisting of bracket repair, rule-based repair, and LLM refinement, ultimately yielding the final Rust project.
Recent work proposes rule-based tools (3; H. Zhang, C. David, Y. Yu, and M. Wang (2023); J. Hong and S. Ryu (2023)), which convert C code to Rust using primarily syntax-based transformations on a terse intermediate representation of compilation. Such translation inherently generates non-idiomatic Rust code with low-level constructs and may also contain semantic inaccuracies. (Pan et al., 2024). Another line of work proposes LLM-based techniques (Eniser et al., 2024; Li et al., 2024; Yang et al., 2024a; Nitin and Ray, 2024) to utilize the code comprehension capability of code LLMs. Due to the lack of parallel C-Rust training data and significant linguistic differences, these approaches often fail to produce semantically equivalent Rust code (Ou et al., 2024).
While automated code translation has been extensively investigated (Jiao et al., 2023; Rozière et al., 2020; Zhu et al., 2022; Yang et al., 2024b; Tao et al., 2024; Dehghan et al., 2025), the translation of C-to-Rust projects poses special challenges:
Challenge-1: Substantial linguistic discrepancies for ensuring safety. C and Rust differ fundamentally in their safety models. C permits loose type checking, unrestricted pointer arithmetic, and manual memory management, whereas Rust enforces strict compile-time type safety, ownership-based memory guarantees, and disciplined reference usage through its borrowing rules. Existing rule-based translators often compromise safety by relying on raw pointers, unsafe blocks, or external C functions (Pan et al., 2024). While LLM-based approaches can generate code with improved safety, they often lack sufficient context understanding (e.g., variable scope or pointer lifetimes), potentially leading to logical errors or memory safety violations in the translated Rust code (Sim et al., 2025). As a result, neither rule-based nor LLM-based methods alone reliably meet Rust’s stringent safety requirements.
Challenge-2: Project-level code dependence. Translating complete C projects poses greater challenges than translating isolated functions because the translated Rust project must preserve cross-module dependencies and consistent project-level structure (Shiraishi and Shinagawa, 2024; Shetty et al., 2024). Code LLMs often struggle with such large-scale context (Yun et al., 2024), leading to broken references, inconsistent APIs, or incorrect module boundaries (Pan et al., 2024; Wang et al., 2025). Consequently, existing approaches tend to perform well only on small-scale programs (e.g., $\leq$100 LoC) (Emre et al., 2021; Eniser et al., 2024), and fall short in real-world project settings.
@@
We evaluate EvoC2Rust using an open-source benchmark Vivo-Bench (Vivo, 2025) and a self-collected dataset of six industrial projects. Experimental results demonstrate that EvoC2Rust significantly outperforms existing baselines in project-level C-to-Rust translation. In comparison with LLM-based approaches, EvoC2Rust improves the compilation pass rate by 17.24–50.84% and the line acceptance rate by 14.32–44.79%, while also achieving marginally better code safety. When compared to rule-based methods, EvoC2Rust attains a 43.59–96.79% higher code safety rate, while maintaining competitive performance in terms of compilation and line acceptance rates. Furthermore, at the module level, EvoC2Rust achieves a 92.25% compilation success rate and an 89.53% test pass rate on industrial projects.
@@
2. Approach
2.1. Project Skeleton Construction
Given a C project, EvoC2Rust first parses the source code using Tree-sitter (41) to extract key syntactic elements and abstract them into a unified project metadata representation. This metadata captures structural information including filenames, include statements, macro definitions, type declarations, global identifiers, function signatures, and a mapping table between identifier declarations to source files. They are then systematically transformed into Rust equivalents via rule-based conversion. For instance, C include dependencies are transformed to Rust use imports, and identifier declarations are converted into Rust pub use statements (e.g., pub use {Rust_filename}::{identifier_name}) using the declaration-to-file table.
Using the converted metadata, EvoC2Rust constructs a Rust project skeleton—comprising Rust constructs automatically translated from C definitions, macros, and function signatures—that accurately preserves the structural organization of the original C project. Each C function is assigned a type-checked dummy counterpart in Rust using its translated signature and the unimplemented!() macro. These placeholders serve as safe stand-ins, allowing the project to compile successfully even before the actual function bodies are translated.
@@ 2.2. Module Translation with Safety-preserving Mapping
Pattern Name: Variadic Argument Mapping   Motivation: Rust does not support C-style variadic functions (e.g., va_list).   Consequence: Current LLMs cannot directly translate C’s variadic features.   Solution: Define a VaList type and va_format! macro: pub type VaList<’a> = &’a [&’a dyn Display];
macro_rules! va_format { … }
  Example: C Code: int snprintf(char *str, size_t size, const char *format, …)
Rust Code: pub fn snprintf(mut buf: Ptr<u8>, size: usize, format: Ptr<u8>,
   va: VaList) -> i32 {
    let mut fmt = format.to_string(); …
    return length; }
@@ 2.3. Post-Generation Repair
Table 3. Statistics of the Datasets
[…]
C2R-Bench avl 9 836 29 29 9 121
bzp 18 2258 92 82 14 17
md5 2 324 8 20 1 37
sha256 2 280 9 11 2 12
rapidlz 7 748 28 38 10 12
cmptlz 25 3724 122 166 27 23
Total 63 8170 288 346 63 222
@@ 3.1. Comparison Methods
We compare EvoC2Rust against three categories of C-to-Rust translation methods: rule-based (C2Rust), LLM-based (Self-Repair and Tymcrat), and hybrid techniques (C2SaferRust). We also incorporate a direct LLM prompting baseline that translates code without repository context. Specifically, the baseline methods evaluated in our study are as follows:
  • C2Rust (3): a rule-based C-to-Rust translator that converts C code to Rust via AST analysis and manually defined transformation rules.
  • C2SaferRust (Nitin et al., 2025): a hybrid technique built on C2Rust, augmented with LLM-based post-processing and test-driven optimization to enhance the safety and idiomaticity of the translated Rust code. We follow the original configuration and use 5 optimization iterations.
  • Self-Repair (Khatry et al., 2025): a multi-agent self-repair framework for project-level C-to-Rust translation. It incrementally performs file translation and uses iterative compilation and testing feedback to correct errors automatically. We adopt the default setting of 3 self-repair rounds per task.
  • Tymcrat (Hong and Ryu, 2025b): a project-level translation method that improves translation quality through type migration. It generates multiple candidate Rust signatures and refines them iteratively using compiler feedback. We use the default configuration, which generates 4 candidate signatures per function.
@@ 3.2. Datasets
We evaluate EvoC2Rust using two benchmarks: the open-source Vivo-Bench (Vivo, 2025), and our newly introduced C2R-Bench, which comprises six industrial projects.
Vivo-Bench is a collection of 19 algorithmic C projects from the 2025 Vivo C-to-Rust Innovation Competition, each containing 1–3 files (33–630 LoC per file). Two senior engineers produced the reference Rust translations with LLM assistance. Subsequently, they validated the translations by adding tests to achieve complete function coverage, resulting in a total of 113 top-level test cases.
@@
Each project in both Vivo-Bench (Vivo, 2025) and C2R-Bench provides four artifacts: the original C source code, C test cases, the reference Rust translation, and corresponding Rust test cases. Table 3 summarizes the dataset statistics.
@@ 3.3. Evaluation Strategy and Metrics
  • Incremental Compilation Pass Rate (ICompRate): The proportion of translated modules that compile successfully when integrated step-wise into the project skeleton, measuring syntactic correctness  (Shiraishi and Shinagawa, 2024).
  • Line Acceptance Rate (AccRate) (Ziegler et al., 2022): This metric evaluates the fidelity of the initial translation by comparing it to the manually corrected version. Precision captures the percentage of correct lines in the initial output, while Recall measures their preservation in the final code.
  • Code Safety Rate (SafeRate) (Ling et al., 2022): The proportion of generated Rust code that is memory-safe, i.e., free of unsafe functions or blocks.
@@
  • Test Pass Rate (TestRate) (Khatry et al., 2025): The percentage of generated modules that pass all associated unit tests.
@@ 3.4. Implementation Details
C2Rust (3) 100 100 100 0
C2SaferRust (Nitin et al., 2025) 100 100 100 60.00
@@
Self-Repair (Khatry et al., 2025) 84.21 92.08 89.24 87.60
Tymcrat (Hong and Ryu, 2025b) 87.35 92.11 92.44 97.88
@@
Self-Repair (Khatry et al., 2025) 52.63 33.12 38.36 77.57
Tymcrat (Hong and Ryu, 2025b) 81.76 76.82 74.47 87.11
@@
C2Rust (3) 99.28 98.99 98.98 1.83
C2SaferRust (Nitin et al., 2025) 99.28 97.47 97.68 48.24
@@
Self-Repair (Khatry et al., 2025) 49.21 72.35 13.79 83.92
Tymcrat (Hong and Ryu, 2025b) 72.02 77.82 74.94 95.92
@@
Self-Repair (Khatry et al., 2025) 39.40 45.38 10.65 87.15
Tymcrat (Hong and Ryu, 2025b) 70.56 66.72 65.91 82.75
@@
We implement EvoC2Rust using DeepSeek-V3 (DeepSeek-AI, 2025) and Qwen3-32B (Alibaba, 2025) as the foundation models. Translation tasks use greedy decoding with max_tokens set to 4096. We leverage Tree-sitter v0.22.3 (41) for C code parsing, the standard rustc (40) for compilation, and the BGE-M3 model (of Artificial Intelligence, 2025) to generate embeddings and retrieve the top-10 most relevant patterns based on cosine similarity. Iteration limits are enforced to ensure practicality: 5 rounds for bracket repair and 3 rounds for LLM refinement.
We established verified ground-truth implementations to evaluate the line acceptance rate metric. For each project, we first translated the code using the target methods. The resulting Rust artifacts were then iteratively compiled and tested. Any compilation or test failures were manually repaired by three senior engineers with assistance from Claude Sonnet 4 (Anthropic, 2024) until all tests passed.
@@ 4.1. Main Results (RQ1)
Specifically, EvoC2Rust improves compilation success by 44.63% and line acceptance by 54.38% over Self-Repair on C2R-Bench with DeepSeek-V3 (42.95%/55.66% with Qwen3-32B). Self-Repair’s file-level processing often exceeds LLM context limits, leading to code truncation. Compared to Tymcrat, EvoC2Rust achieves a 21.82% higher compilation rate on C2R-Bench with DeepSeek-V3 (11.79% with Qwen3-32B). While Tymcrat employs diverse signature generation and context enrichment to reduce initial errors, its lack of expert-derived structured translation guidance ultimately constrains its effectiveness in complex scenarios. Direct Prompting performs poorest, with a 38.42% average compilation success and low line acceptance (39.13% precision, 40.64% recall). This is primarily due to the inherent linguistic gaps between C and Rust.